# MRHaD: Mixed Reality-based Hand-Drawn Map Editing Interface for Mobile Robot Navigation

### v1

@@ II-A Robot Navigation Interfaces
Traditional navigation interfaces have primarily used 2D graphical user interfaces (GUIs) to display sensor data, such as camera, sonar, and LiDAR readings, and to allow users to specify goals or plan paths. Previous work introduced systems that overlay 2D maps with sensor-based environment information, improving situational awareness for human operators [7, 8]. Subsequent studies explored more intuitive interactions, such as varying methods for setting destinations [9] and block-based approaches for control and path specification [10]. In the context of ROS, RViz remains a popular tool, providing both 2D and 3D visualization of the robot’s environment [11].
@@
Map Editting Accuracy $\checkmark$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$
Editing time required $\times$ $\times$ $\checkmark$ $\times$ $\checkmark$
@@
Figure 3: System Structure and Data Flow. Once the drawing operation is completed, the vertex coordinates and identification number (HRZ-ID) of the HRZ are sent to ROS 2. ROS 2 converts the received coordinates appropriately, stores them along with HRZ-ID, and then adds the HRZ to the map received from Kachaka. When HRZ is deleted, the corresponding HRZ-ID is sent from the HoloLens 2 to ROS 2, which then removes only the specified HRZ from the map.
@@ III SYSTEM DESIGN
The structure and data flow of the proposed system are shown in Fig. 3. The system consists of a user, HoloLens 2, the Robot Operating System 2 (ROS 2), and the mobile robot, Kachaka. The newly developed components in this paper are shown in green.
The user wears the HoloLens 2 and draws Hand-drawn Restricted Zones (HRZ) using hand gestures. The drawn HRZ is visualized as a green polygon on the HoloLens 2. In ROS 2, the addition of HRZ is realized by connecting the received vertex coordinates to form a polygon and marking its edges and interior cells as occupied. Deletion is achieved by resetting the map and reapplying all HRZ except the one marked for deletion. HRZ-ID and vertex coordinates of each HRZ are stored in an internal database of the HoloLens 2, allowing the system to restore them upon restart. Fig. 2 shows an example of HRZ displayed on the HoloLens 2 and its reflection on the ROS 2 cost map.
@@ IV-C Coordinate Data Management
6:        SaveDatabase($database$, $hrz\_id$, $coordinates$,       $rotation$)
7:        OverwriteJsonFile($database$)
@@ V-C2 Map Reflection Accuracy
Stage/ Obs. Accuracy Precision Recall Specificity F-1 Score

### v2

@@ II-A Robot Navigation Interfaces
Traditional navigation interfaces have primarily used 2D graphical user interfaces (GUIs) to display sensor data, such as camera, sonar, and LiDAR readings, and to allow users to specify goals or plan paths. Previous work introduced systems that overlay 2D maps with sensor-based environment information, improving situational awareness for human operators [7, 8]. Subsequent studies explored more intuitive interactions, such as varying methods for setting destinations [9] and block-based approaches for control and path specification [10]. Furthermore, RViz remains a popular tool, providing both 2D and 3D visualization of the robot’s environment [11].
@@
Map Editing Accuracy $\checkmark$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$
Required Editing Time $\times$ $\times$ $\checkmark$ $\times$ $\checkmark$
@@
Figure 3: System Structure and Data Flow. Once the drawing operation is completed, the vertex coordinates and identification number (HRZ-ID) of the HRZ are sent to ROS 2. ROS 2 converts the received coordinates appropriately, stores them along with HRZ-ID, and then adds the HRZ to the map received from Kachaka. When HRZ is deleted, the corresponding HRZ-ID is sent from HoloLens 2 to ROS 2, which then removes only the specified HRZ from the map.
@@ III SYSTEM DESIGN
The user wears HoloLens 2 and draws Hand-drawn Restricted Zones (HRZ) using hand gestures. The drawn HRZ is visualized as a green polygon on HoloLens 2. Fig. 2 shows an example of HRZ displayed on HoloLens 2 and its reflection on the cost map within the Robot Operating System 2 (ROS 2).
In ROS 2, the addition of HRZ is realized by connecting the received vertex coordinates to form a polygon and marking its edges and interior cells as occupied. Deletion is achieved by resetting the map and reapplying all HRZ except the one marked for deletion. HRZ-ID and vertex coordinates of each HRZ are stored in an internal database of HoloLens 2, allowing the system to restore them upon restart.
The structure and data flow of the proposed system are shown in Fig. 3. The system consists of a user, HoloLens 2, ROS 2, and the mobile robot, Kachaka. The newly developed components in this paper are shown in green.
@@ IV-C Coordinate Data Management
[…]
@@ V-C2 Map Reflection Accuracy
Stage/
Obs.
Accuracy Precision Recall Specificity F-1 Score